# Cognitive Control Architecture (CCA): A Lifecycle Supervision Framework for Robustly Aligned AI Agents

### v1

@@ 1 Introduction
Large Language Model (LLM) agents are increasingly deployed in autonomous systems, capable of accomplishing complex real-world tasks through extensive tool usage (xi2025rise; deng2023mind2web). However, this growing autonomy introduces critical security challenges: their limited reasoning capabilities, particularly their failure to consistently recognize high-risk situations (zhang2024agent), significantly expands the attack surface. Among numerous threats, Indirect Prompt Injection (IPI) is particularly concerning. IPI attacks hijack agent behavior by subtly contaminating external information sources (greshake2023not; debenedetti2024agentdojo; naihin2023testing; zhan2024injecagent), exploiting the fundamental vulnerability where agents struggle to distinguish between ‘instructions’ and ‘data’ within processing contexts. This allows attackers to hijack the agent’s tool usage, forcing agents to execute unauthorized operations and deviate from their intended purpose (owasp2025top10). Existing defense mechanisms are caught in a fundamental trade-off between security and functionality, since overly strict safety rules often limit what the agent can actually do, while greater autonomy expands the attack surface. The fragmented nature of these defenses prevents end-to-end integrity assurance, leaving them ill-equipped to counter sophisticated Indirect Prompt Injection (IPI) attacks. Such attacks succeed by inducing malicious operations, such as data exfiltration, that are crafted to appear compliant with local task flows. This exposes the inability of current paradigms to resolve the multi-dimensional trade-offs between security, functionality, and efficiency against these threats (yan2025attack; yang2025drunkagent; jiang2025policy; eghtesad2023multi).
@@
Current SOTA defense mechanisms are still largely fragmented and static, resulting in a difficult compromise between security, functionality, and efficiency. Runtime checks (zhu2025melon; jia2024task) focus on isolated actions, making them blind to long-range malicious plans. Architectural defenses (wu2024isolategpt; wu2024system) impose significant performance overhead or functional limitations. Meanwhile, training-time approaches (wallace2024instruction; chen2025struq) struggle to generalize against unseen or optimization-based attacks (zou2023universal). Consequently, these single-point solutions are inherently ill-equipped to detect sophisticated deceptively aligned agents, where a deceptive agent’s behavior appears locally coherent while globally deviating from user intent. This phenomenon is not a corner case but rather a scalability challenge that inevitably emerges as the reasoning and planning capabilities of models advance.
@@
To address this challenge, this paper introduces the Cognitive Control Architecture (CCA), a principled framework designed to elevate agent security from isolated, post-hoc behavioral validation to continuous, full-lifecycle cognitive supervision. CCA provides systemic safety through two synergistic pillars. First, an ”Intent Graph” establishes both control-flow and data-flow integrity by pre-generating a baseline of legitimate cognitive flow, a concept that reinforces the plan-and-execute paradigms (yao2023react) to efficiently thwart overt planning deviations. Second, upon detecting a deviation, an innovative Tiered Adjudicator intervenes, performing deep causal reasoning via a multifaceted ”Intent Alignment Score”.
@@ 2 Background and Related Work
Large Language Model (LLM) agents are vulnerable to Indirect Prompt Injections (IPIs), which exploit their inability to distinguish between trusted instructions and untrusted data. Attackers embed malicious instructions within external data sources to hijack the agent’s behavior and deviate it from the user’s intent, as illustrated in Figure 1. These attacks have evolved from early, generic techniques that exploit general LLM weaknesses, such as context ignoring (schulhoff2023ignore) or false completion (liu2024formalizing), into more sophisticated, context-specific penetrations. Modern variants include polluting web content to attack web agents (greshake2023not) or manipulating file system metadata to attack computer-controlling agents (su2025survey). This trend towards context-dependent attacks makes malicious behavior far more covert, as it targets deep vulnerabilities in the agent’s cognitive processes by exploiting trust in specific environmental interaction patterns (yan2025attack; yang2025drunkagent; jiang2025policy; eghtesad2023multi).
Research into agent security has revealed that vulnerability to Indirect Prompt Injection (IPI) attacks stems from two core deficits: a lack of coherent, long-term task planning, and an inherent lack of risk awareness when processing external information. Existing defense paradigms tend to address only one of these, leaving critical blindspots. Runtime verification, used by systems like MELON (zhu2025melon), validates actions based on their causal independence from user intent but suffers from a ”causal blindspot,” struggling to detect conditional attacks where malicious logic is intertwined with legitimate task components. System-level isolation frameworks like IsolateGPT (wu2024isolategpt) use sandboxing but often incur significant performance overhead and fail against attacks that abuse legitimate capabilities via semantic manipulation. Training-time alignment (ouyang2022training) attempts to instill intrinsic robustness but suffers from poor generalization against unseen or optimization-based attacks like GCG (zou2023universal). Finally, Control-Flow Integrity (CFI) approaches, such as IPIGuard (an2025ipiguard), correctly address the long-term planning deficit with tool-dependency graphs. However, by primarily enforcing topological correctness, they fail to address the agent’s lack of risk awareness. Their dynamic mechanisms still rely on a naive LLM core, allowing attackers to exploit the causal blindspot by manipulating the parameters of a legitimate tool call. In summary, existing paradigms are ill-equipped to handle the growing complexity of IPI attacks, which motivates our proposal of the Cognitive Control Architecture (CCA).
@@ 3.3 Pillar II: High-Precision Anomaly Adjudication
Semantic Alignment ($S_{\text{sem}}$): This score quantifies the semantic relevance of the deviated action to the user’s original goal $\mathcal{G}_{u}$. It is computed using the all-MiniLM-L6-v2(reimers2021allminlm) model to obtain 384-dimensional dense vector embeddings of the natural language description of the deviated action $a_{t}$ and its accompanying justification, and the user’s goal $\mathcal{G}_{u}$. The score is then calculated as the cosine similarity between these embeddings (Equation 2):
@@
Consistent with agent paradigms like ReAct (yao2023react), where actions are conditioned on the most recent observation, the immediate output of the preceding tool call is the primary attack surface for IPI. Our $S_{\text{prov}}$ score is therefore designed to precisely evaluate the reliability of the external information source (e.g., API, URL) that triggered the deviation. The trust score $T(\text{source})$ is initialized with a preset value and dynamically updated after each interaction based on a reinforcement learning-like rule (Equation 4):
@@ 4.1 Experimental Setup
Evaluation Benchmark. We evaluate CCA on AgentDojo (debenedetti2024agentdojo), a dynamic, stateful benchmark with 97 multi-turn tasks across domains like email, banking, and workspaces. It uses deterministic code evaluation for reliability, ideal for testing CCA’s global supervision. [^1] Details in (debenedetti2024agentdojo); access at https://agentdojo.spylab.ai.
Models. We use DeepSeek-V3.1 (liu2024deepseek) and Kimi K2 (team2025kimi) as core agents, consistent for planning and execution. Baselines match the agent model. The Adjudicator $M_{\text{adj}}$ is fixed as DeepSeek-V3.1.
Attacks. We test against four AgentDojo IPI attacks: Direct (debenedetti2024agentdojo), Ignore Prev (rahman2025fine; schulhoff2023ignore), System Msg (debenedetti2024agentdojo), and Important Msgs (debenedetti2024agentdojo), covering simple to complex strategies.
Baselines. We compare CCA against four baseline methods representing different defensive paradigms: prompt engineering-based defenses Repeat Prompt and Spotlight (hines2024defending); an external model-based DeBERTa classifier (ProtectAI_Deberta_PI_2024); and a state-of-the-art (SOTA) runtime verification method, MELON. Additionally, we report results for a No Defense agent as a reference.

### v2

@@ 1 Introduction
Large Language Model (LLM) agents are increasingly deployed in autonomous systems, capable of accomplishing complex real-world tasks through extensive tool usage (Xi et al., 2025; Deng et al., 2023). However, this growing autonomy introduces critical security challenges: their limited reasoning capabilities, particularly their failure to consistently recognize high-risk situations (Zhang et al., 2024), significantly expands the attack surface. Among numerous threats, Indirect Prompt Injection (IPI) is particularly concerning. IPI attacks hijack agent behavior by subtly contaminating external information sources (Greshake et al., 2023; Debenedetti et al., 2024; Naihin et al., 2023; Zhan et al., 2024), exploiting the fundamental vulnerability where agents struggle to distinguish between ‘instructions’ and ‘data’ within processing contexts. This allows attackers to hijack the agent’s tool usage, forcing agents to execute unauthorized operations and deviate from their intended purpose (OWASP Foundation, 2023). Existing defense mechanisms are caught in a fundamental trade-off between security and functionality, since overly strict safety rules often limit what the agent can actually do, while greater autonomy expands the attack surface. The fragmented nature of these defenses prevents end-to-end integrity assurance, leaving them ill-equipped to counter sophisticated Indirect Prompt Injection (IPI) attacks. Such attacks succeed by inducing malicious operations, such as data exfiltration, that are crafted to appear compliant with local task flows. This exposes the inability of current paradigms to resolve the multi-dimensional trade-offs between security, functionality, and efficiency against these threats (Yan et al., 2025; Yang et al., 2025; Jiang et al., 2025; Eghtesad et al., 2023).
@@
Current SOTA defense mechanisms are still largely fragmented and static, resulting in a difficult compromise between security, functionality, and efficiency. Runtime checks (Zhu et al., 2025; Jia et al., 2024) focus on isolated actions, making them blind to long-range malicious plans. Architectural defenses (Wu et al., 2024b; a) impose significant performance overhead or functional limitations. Meanwhile, training-time approaches (Wallace et al., 2024; Chen et al., 2025) struggle to generalize against unseen or optimization-based attacks (Zou et al., 2023). Consequently, these single-point solutions are inherently ill-equipped to detect sophisticated deceptively aligned agents, where a deceptive agent’s behavior appears locally coherent while globally deviating from user intent. This phenomenon is not a corner case but rather a scalability challenge that inevitably emerges as the reasoning and planning capabilities of models advance.
@@
To address this challenge, this paper introduces the Cognitive Control Architecture (CCA), a principled framework designed to elevate agent security from isolated, post-hoc behavioral validation to continuous, full-lifecycle cognitive supervision. CCA provides systemic safety through two synergistic pillars. First, an ”Intent Graph” establishes both control-flow and data-flow integrity by pre-generating a baseline of legitimate cognitive flow, a concept that reinforces the plan-and-execute paradigms (Yao et al., 2023) to efficiently thwart overt planning deviations. Second, upon detecting a deviation, an innovative Tiered Adjudicator intervenes, performing deep causal reasoning via a multifaceted ”Intent Alignment Score”.
@@ 2 Background and Related Work
Large Language Model (LLM) agents are vulnerable to Indirect Prompt Injections (IPIs), which exploit their inability to distinguish between trusted instructions and untrusted data. Attackers embed malicious instructions within external data sources to hijack the agent’s behavior and deviate it from the user’s intent, as illustrated in Figure 1. These attacks have evolved from early, generic techniques that exploit general LLM weaknesses, such as context ignoring (Schulhoff et al., 2023) or false completion (Liu et al., 2024b), into more sophisticated, context-specific penetrations. Modern variants include polluting web content to attack web agents (Greshake et al., 2023) or manipulating file system metadata to attack computer-controlling agents (Su et al., 2025). This trend towards context-dependent attacks makes malicious behavior far more covert, as it targets deep vulnerabilities in the agent’s cognitive processes by exploiting trust in specific environmental interaction patterns (Yan et al., 2025; Yang et al., 2025; Jiang et al., 2025; Eghtesad et al., 2023).
Research into agent security has revealed that vulnerability to Indirect Prompt Injection (IPI) attacks stems from two core deficits: a lack of coherent, long-term task planning, and an inherent lack of risk awareness when processing external information. Existing defense paradigms tend to address only one of these, leaving critical blindspots. Runtime verification, used by systems like MELON (Zhu et al., 2025), validates actions based on their causal independence from user intent but suffers from a ”causal blindspot,” struggling to detect conditional attacks where malicious logic is intertwined with legitimate task components. System-level isolation frameworks like IsolateGPT (Wu et al., 2024b) use sandboxing but often incur significant performance overhead and fail against attacks that abuse legitimate capabilities via semantic manipulation. Training-time alignment (Ouyang et al., 2022) attempts to instill intrinsic robustness but suffers from poor generalization against unseen or optimization-based attacks like GCG (Zou et al., 2023). Finally, Control-Flow Integrity (CFI) approaches, such as IPIGuard (An et al., 2025), correctly address the long-term planning deficit with tool-dependency graphs. However, by primarily enforcing topological correctness, they fail to address the agent’s lack of risk awareness. Their dynamic mechanisms still rely on a naive LLM core, allowing attackers to exploit the causal blindspot by manipulating the parameters of a legitimate tool call. In summary, existing paradigms are ill-equipped to handle the growing complexity of IPI attacks, which motivates our proposal of the Cognitive Control Architecture (CCA).
@@ 3.3 Pillar II: High-Precision Anomaly Adjudication
Semantic Alignment ($S_{\text{sem}}$): This score quantifies the semantic relevance of the deviated action to the user’s original goal $\mathcal{G}_{u}$. It is computed using the all-MiniLM-L6-v2(Reimers, 2021) model to obtain 384-dimensional dense vector embeddings of the natural language description of the deviated action $a_{t}$ and its accompanying justification, and the user’s goal $\mathcal{G}_{u}$. The score is then calculated as the cosine similarity between these embeddings (Equation 2):
@@
Consistent with agent paradigms like ReAct (Yao et al., 2023), where actions are conditioned on the most recent observation, the immediate output of the preceding tool call is the primary attack surface for IPI. Our $S_{\text{prov}}$ score is therefore designed to precisely evaluate the reliability of the external information source (e.g., API, URL) that triggered the deviation. The trust score $T(\text{source})$ is initialized with a preset value and dynamically updated after each interaction based on a reinforcement learning-like rule (Equation 4):
@@ 4.1 Experimental Setup
Evaluation Benchmark. We evaluate CCA on AgentDojo (Debenedetti et al., 2024), a dynamic, stateful benchmark with 97 multi-turn tasks across domains like email, banking, and workspaces. It uses deterministic code evaluation for reliability, ideal for testing CCA’s global supervision. [^1] Details in (Debenedetti et al., 2024); access at https://agentdojo.spylab.ai.
Models. We use DeepSeek-V3.1 (Liu et al., 2024a) and Kimi K2 (Team et al., 2025) as core agents, consistent for planning and execution. Baselines match the agent model. The Adjudicator $M_{\text{adj}}$ is fixed as DeepSeek-V3.1.
Attacks. We test against four AgentDojo IPI attacks: Direct (Debenedetti et al., 2024), Ignore Prev (Rahman et al., 2025; Schulhoff et al., 2023), System Msg (Debenedetti et al., 2024), and Important Msgs (Debenedetti et al., 2024), covering simple to complex strategies.
Baselines. We compare CCA against four baseline methods representing different defensive paradigms: prompt engineering-based defenses Repeat Prompt and Spotlight (Hines et al., 2024); an external model-based DeBERTa classifier (ProtectAI, 2024); and a state-of-the-art (SOTA) runtime verification method, MELON. Additionally, we report results for a No Defense agent as a reference.